OpenStudy (anonymous):

Which of the following are quadratic functions? Check all that apply. A. y=5+2x B. y=2(x-3x^2+1) C. y=1/2x^2-13x+4 D. y=x/x-1

OpenStudy (anonymous):

@KeithAfasCalcLover do you mind helping with this?

OpenStudy (anonymous):

This is the quadratic formula - x = [ -b ± sqrt(b^2 - 4ac) ] / 2a

OpenStudy (anonymous):

Sure Ali :) Im just going to make 'em a little neat: We have: \[\eqalign{ &A.\phantom{a}y=5+2x \\ &B.\phantom{a}y=2(x-3x^2+1) \\ &C.\phantom{a}y=\frac{x^2}{2}-13x+4 \\ &D.\phantom{a}y=\frac{x}{x-1} \\ }\] Right?

OpenStudy (anonymous):

yes sir! :)

OpenStudy (anonymous):

Beautiful. So then the criteria for a quadratic I would say...is: -It has to be a second degree polynomial (meaning it has to have the "\(x^2\)" part lol -And it has to be re-arrangeable to the form \(ax^2+bx+c\). Where \(\{a,b,c\in R|a,b,c\neq0\}\). That means that the co-efficients have to be regular numbers like \(5.344\) or \(-6\frac{3}{4}\) and cant be zero. So lets look at the first one: Well the co-efficients are real numbers buuuut...it's missing the "\(x^2\)" part right?

OpenStudy (anonymous):

Correct, so A would be out of the picture as well as D?

OpenStudy (anonymous):

Nice job Ali. Exactly. So apparently, nothing is wrong with C and D. They both have an "\(x^2\)" and their \(a,b,\) and \(c\) are all real numbers. So I guess that's it lol

OpenStudy (anonymous):

So it's C and D, not B and C?

OpenStudy (anonymous):

Ehh, well its B and C remember? Cause D is not valid

OpenStudy (anonymous):

That's what I thought but earlier you had put "nothing is wrong with C and D" so I got confused :o